Find the area of the shaded region. The graph depicts the standard normal distribution with mean 0 and standard deviation 1. -2.36-1.77-1.18-0.59 0.59 1.18 1.77 2.36 2 O0.7224 0.2776 0.2224 0.2190
Answer #1

Solution:

Given that

P ( Z > 0.59 )

Using standard normal table

= 1 - P ( Z < 0.59 ) = 0.7224

P ( Z < 0.59 ) = 1 - 0.7224

                     = 0.2776

The area of shaded region is 0.2776

P ( Z > 0.59 ) = 2776

Option 0.2776 is correct.